# Robotic Sorting Systems: Robot Management and Layout Design Optimization

### v1

@@ Abstract
In the contemporary logistics industry, automation plays a pivotal role in enhancing production efficiency and expanding industrial scale. Autonomous mobile robots, in particular, have become integral to the modernization efforts in warehouses. One noteworthy application in robotic warehousing is the robotic sorting system (RSS), distinguished by its characteristics such as cost-effectiveness, simplicity, scalability, and adaptable throughput control. While previous research has focused on analyzing the efficiency of RSS, it often assumed an ideal robot management system ignoring potential queuing delays by assuming constant travel times. This study relaxes this assumption and explores the quantitative relationship between RSS configuration parameters and system throughput. We introduce a novel robot traffic management method, named the rhythmic control for sorting scenario (RC-S), for RSS operations, equipped with an estimation formula establishing the relationship between system performance and configurations. Simulations validate that RC-S reduces average service time by 10.3% compared to the classical cooperative A* algorithm, while also improving throughput and runtime. Based on the performance analysis of RC-S, we further develop a layout optimization model for RSS, considering RSS configuration, desired throughput, and costs, to minimize expenses and determine the best layout. Numerical studies show that at lower throughput levels, facility costs dominate, while at higher throughput levels, labor costs prevail. Additionally, due to traffic efficiency limitations, RSS is well-suited for small-scale operations like end-of-supply-chain distribution centers. Keywords: Logistics; Robotic sorting system; Robot management system; Performance evaluation; Layout design
@@ 1 Introduction
Improved logistics and delivery services have fueled the rapid growth of e-commerce in the 21st century. The development of stable supply chains, same-day or next-day delivery options, and hassle-free return policies increase consumer confidence in online shopping (BusinessWire,, 2023). In order to handle the increasing volume of online orders, a new generation of warehouses specifically catering to individual customers has become a hot topic for logistics companies. This kind of warehouse efficiently meets the demand for small orders with tight delivery schedules through the implementation of automated equipments (Boysen et al., 2019b, ). For instance, robotic mobile fulfill system (RMFS), which employs a rack-moving mechanism, is widely used in intra-warehouse logistics, such as Amazon’s KIVA systems (Wurman et al.,, 2008). The main activities performed in warehouse include: (1) receiving, (2) transfer and put away, (3) order picking/selection, (4) accumulation/sorting, (5) cross-docking, and (6) shipping (De Koster et al.,, 2007).
In this paper, we focus on the sorting process. In a distribution center or warehouse where batching and/or zoning policies are implemented within the order picking process, sorting is required to split the batch and consolidate items per customer order or destination for shipping. Conventional sorting systems commonly employ sorting conveyors equipped with different sorter technologies. The sorters above move steadily, sequentially pass through each outbound station, and release the loaded parcels at suitable locations (Boysen et al., 2019a, ). These systems are highly appreciated for its efficiency and stability, while suffering from the inflexibility and significant space occupation (Boysen et al.,, 2023).
@@
A novel sorting system utilizing mobile robots attracts increasing attention in recent years, which is called the robotic sorting system (RSS) (Zou et al.,, 2021). Figure 1(a) presents a dual-layer implementation of RSS. The site is dotted with considerable number of outlets, each of which is associated with a specific sorting category, gathering all the packages of that category. The hardware for the outlet can be a bin, a cage cart, or a chute on the floor, which connects to the packing equipment on the lower level. The loading stations are located on the periphery of the site, where workers are responsible for loading parcels onto the robots. Sorting robots are the devices responsible for carrying out the sorting operations, serving as the conveyor belts in conventional sorting systems. A top-mounted tray or conveyor device allows the robot to load or unload shipments. Figure 1(b) shows a specific model of sorting robot developed by Geekplus, a global technology company specialized in smart logistics and robots. The operational cycle of a sorting robot is as follows:
@@
The independent and modular nature of robotic operations enables the system design process to become more flexible. During off-hours break, warehouse managers can easily modify the control software to change the positions of loading stations, correspondences between outlets and destinations, and the topology of the aisle network (Xu et al.,, 2022). Compared with conveyor belt equipment, RSS possesses the advantage of scalability and throughput flexibility, which are highly suited to the high-demand and dynamic nature of the current e-commerce environment (Azadeh et al.,, 2017). To promote this advantage, robotics companies introduce an innovative business model called robots-as-a-service, in which robotics companies offer customers the opportunity to rent robots for use in manufacturing. As a result, the customers no longer have to contend with the high costs typically associated with purchasing advanced automation tools outright. In RSS, warehouse manager can adjust the number of robots based on sorting demand fluctuations, thereby reducing operations costs.
@@
However, the flexibility of robot system also brings additional operational challenge, namely the multi-agent path finding (MAPF) problem, which aims to resolve the potential conflicts among the trajectories of robots. The MAPF problem has been proven to be NP-hard (Yu and LaValle,, 2013), which means that finding a relatively good solution can be time-consuming, especially when the scale of the robot system increases, resulting in the instability of RSS efficiency. Shi et al., (2021) analyzed a dataset from the China Post sortation center and found that the congestion effect significantly hampers robot efficiency in an RSS, particularly when dealing with a substantial parcel flow. Existing studies often overlook issues related to robot coordination, making the idealized assumption that there are no conflicts or deadlocks. Additionally, these studies tend to assume a fixed route travel time excluding the potential queuing delays. As a result, their outcomes may become distorted, especially in situations of high traffic load, as the actual travel time of the robots deviates from the free-flow travel time.
@@
Given the existing gap, this paper aims to address critical challenges in designing an efficient RSS, focusing on three key areas. First, we seek to ensure the system operates both stably and efficiently in a setting where multiple robots collaborate, a non-trivial challenge given the complexities of robot interaction. Second, the development of a precise model to quantify congestion effects within the RSS is essential. Such a model will enable accurate throughput estimation, which is vital for understanding the relationship between the size of the facility and its sorting capacity. Third, we emphasize the need for management strategies that exploit the flexibility of RSS to adapt to fluctuating demands cost-effectively. These targeted areas of improvement aim to enhance the overall performance and scalability of RSS implementations.
@@
To address the aforementioned challenges and achieve the system optimum, we first explore the potential of centralized control strategy for sorting robots. To fully utilize the global information of the robot control system, taking into account robot operational patterns, and safety distance regulations, we adopt a network-level control strategy for the online coordination of autonomous vehicles and propose a collision-free spatio-temporal path planning method for RSS. Compared to existing MAPF methods, it does not require high communication frequency and tight latency between the central host and each robot. Additionally, we derive an efficient heuristic algorithm to enhance system scalability. Second, building upon this control strategy, under any given system configurations, we conduct a quantitative assessment of the average travel distance of robots and the maximum throughput of the RSS. We introduce an estimation formula that elucidates the relationship between system performance metrics and configuration settings. Finally, we propose an optimization model aimed at minimizing the initial investment and average operations costs under variable demand. The solution includes layout design and resource allocation recommendations. Experiments are conducted to provide several insights about the RSS, such as the cost structure of the RSS across varying unit price conditions. To the best of our knowledge, this paper is among the first to incorporate the traffic issues of robots during operations into the configuration design of the RSS system. Our intention is to bridge operational and strategic planning in multi-robot coordination contexts, and hence contribute a comprehensive theoretical framework to support decision-making in such complex systems.
@@
The remainder of this paper is organized as follows. In section 2, we provide an extensive review of relevant research on RSS systems and conventional sorting systems, highlighting the contributions of this research. Section 3 offers a detailed problem description. In section 4, we introduce an innovative traffic management framework to coordinate multiple robots in a warehouse setting. Based on the proposed framework, section 5 presents a system throughput estimation formula, which provides quantitative measurement of layout configuration’s impact to system efficiency. In section 6, we demonstrate the superiority of our proposed traffic management framework compared with benchmarks, as well as validate the throughput estimation formula, through comprehensive experiments. Moving on to section 7, we propose a layout optimization model to minimize the total costs and introduce an efficient solution approach. Section 8 presents a sensitivity analysis to investigate the optimal layout design and the corresponding cost structure under different unit cost, and insights distilled from the results. Finally, section 9 provides a summary of the entire paper.
@@ 2 Literature Review
The cost savings in labor through automation are particularly crucial in the picking process, and higher sorting depth enables parcel companies to offer convenient and low-cost same-day delivery services (Dekhne et al.,, 2019). However, the operations research community has scarcely explored the realm of robotized sorting systems, which represent the latest advance of warehouse automation. Among the existing literature, studies mainly focus on small parcel sorting systems in distribution centers. Boysen et al., (2023) conducted research on order sorting system, which involves additional tasks including Piece-to-Order assignment and Order-to-Collection Point assignment. Therefore, the primary focus of the referred study is not solely on scheduling and control of robots, and the travel time of robots to a specific outlet is assumed to be constant.Shi et al., (2021)’s study, on the other hand, investigates a human-robot hybrid sorting system, in which the scale of robot sorting is fixed, while the manual capacity is adjustable according to the demand. Regrettably, in this context, the considerable flexibility inherent in the robotic system has not been fully harnessed.
@@
In addition to the two aforementioned references, the remaining literature on robotic sortation, including this study, reorients the focus back to the original RSS itself. This body of work delves into the comprehensive assessment of transportation efficiency within the robot fleet under various system configurations. In the majority of these studies, robots’ sorting tasks are modeled as services in queuing network models, and the total throughput is the key performance evaluation metrics (Zou and Chen,, 2020; Zi and Gao, 2020a, ; Zi and Gao, 2020b, ; Zou et al.,, 2021; Xu et al.,, 2022). Zi and Gao, 2020a ; Zi and Gao, 2020b assumed that each robot moves independently. Zou and Chen, (2020) introduced a semi-open queuing network (SOQN) model and subsequently improved its theoretical aspects in their follow-up research (Zou et al.,, 2021). The SOQN models in these two studies take into account the queuing behavior of robots at loading stations and outlets, partially addressing the issue of robot occupancy on the aisle. Furthermore, their second study compares how the minimum total costs of the system varied under different network topology and different unit cost of system components, given the target sorting efficiency. Xu et al., (2022) further investigated an RSS with a parcel-to-loading-station assignment mechanism, which could be regarded as a pre-sorting strategy. To simplify the model, the study removes the constraint on the number of robots, resulting in an open queuing network (OQN) model. Experimental results demonstrates that the introduction of pre-sorting effectively reduces the average travel distance of the robots but might lead to extra congestion upstream at the same time. Additionally, Liu et al., 2019b and Tan et al., (2021) focused on the task assignment problem with the objective of minimizing the sorting makespan. Both adopts a travel time model (closed-form travel time expression, defined by Azadeh et al., (2017)), and the second research formulates a mixed-integer programming model. Compared to queuing network models, travel time models exhibit significant distortion when the number of robots is high, as they allow multiple robots to simultaneously occupy a facility.
@@
Throughout the existing research, a limited number of studies consider the floor space cost while others relax the constraints on aisle resources. Furthermore, finding collision-free paths for multiple vehicles has been proved to be an NP-hard problem (Surynek,, 2010), and some state-of-the-art algorithms (such as the CROTPN-based model in Sun et al.,, 2018) also struggle to completely eliminate the issue of road congestion resulting from an increasing number of robots. However, except the study by Shi et al., (2021), which employs a fitting method to predict the positive correlation between traffic flow and congestion, most of the studies neglect the mutual influence among multiple robots in aisle occupation for a compact system, leading to an overestimation of aisle capacity. Table 1 presents a comprehensive overview of the literature reviewed on the RSS, highlighting its relevance to our paper.
Next, we review the performance analysis and layout design of sorting systems based on conveyors. The study by Russell and Meller, 2003a comprehensively investigates the layout design problem in an order fulfillment system. Their descriptive model analyzes the impact of the number of workers and workstations on the system’s capacity, considering the costs of labor and equipment, as well as providing appropriate system configuration based on target throughput requirement. The study also validates their analytical results through simulation experiments and concluded the advantages of conveyor belt sorting compared to manual sorting. Bartholdi and Gue, (2000) studied the scenario of a cross-docking terminal, where freight needs to undergo sorting on a dragline conveyor system from strip doors to stack doors. They accounted for three types of congestion in the system and employed a simulated annealing method to examine sorting efficiency and labor costs across various layouts. Besides, the postal service industry faces greater challenges, including higher volume of shipments and tighter time window. Fedtke and Boysen, (2017) examined how design alternatives within the distribution center affect the performance of sorting system. The authors devised an integer programming model to optimize the minimum service distance under various configurations. Additionally, they developed heuristic algorithms for solutions. Similar methodologies have been applied in several other researches (Briskorn et al.,, 2017; Zenker and Boysen,, 2018). Readers are recommended to refer to Boysen et al., 2019a for an in-depth understanding of sortation conveyor systems.
@@
In addition to sortation systems, there are a sufficient number of studies on RMFS system in modern order picking process, which also consider the multi-robot management problem. An important operational metric of order picking is the proportion of effective working time for workers, and thus, related studies often utilize queuing network models to characterize the uncertainty in RMFS. Lamballais et al., (2017) proposed a SOQN model and provided insights into how performance is affected by factors such as the length-to-width ratio, workstation locations, number of robots. This modeling approach effectively captures the system bottleneck associated with robot blocking at workstations, which is validated through simulation. This modeling approach has proven instrumental in addressing challenges related to battery recovery, robot assignment, inventory allocation, and multi-deep layout (Zou et al.,, 2018; Roy et al.,, 2019; Tessensohn et al.,, 2020; Yang et al.,, 2022). Certain studies have relaxed the constraint on the number of robots, depending on the focus of their research (open queuing network: Yuan and Gong,, 2017, closed queuing network: Nigam et al.,, 2014; Roy et al.,, 2019). However, some researchers have recognized the limitations of queuing models in capturing the impact of aisle design on system throughput, and have chosen to address the problem from the perspective of MAPF problem in their studies. Wagner and Choset, (2011) proposed a novel M* algorithm that allowed for the planning of a larger number of robot trajectories compared to the original A* algorithm. However, they faced challenges due to significant memory requirements and time complexity, and could not guarantee a solution. Some researchers designed priority-based algorithms and integrated task assignment to reduce waiting time (Nguyen et al.,, 2019; Liu et al., 2019a, ), sacrificing algorithm optimality in exchange for stability and computational efficiency. Li et al., (2021) introduced an RHCR algorithm framework to tackle the windowed MAPF problem, providing an efficient solution for managing newly arrived orders in warehouse operations. The algorithm demonstrates scalability up to 1,000 agents and yields high-throughput solutions. However, despite their effectiveness, MAPF-based studies have encountered challenges in deriving a closed-form expression for throughput estimation, a critical element for robust system configurations. In summary, there are few studies on RMFS that comprehensively address both operational robot management problems and strategic warehouse design, with considering their interplay. We aim for the findings and framework of this study not only to be applicable to the RSS system, but also to be extendable to these multi-robot systems in future studies.

### v2

@@ Abstract
In the contemporary logistics industry, automation plays a pivotal role in enhancing production efficiency and expanding industrial scale. In particular, autonomous mobile robots have become integral to modernization efforts in warehouses. One noteworthy application in robotic warehousing is the robotic sorting system (RSS), which is distinguished by its cost-effectiveness, simplicity, scalability, and adaptable throughput control. Previous research on RSS efficiency often assumed an ideal robot management system, ignoring potential traffic delays and assuming constant travel times. We introduce a novel robot traffic management method, named Rhythmic Control for Sorting Scenario (RC-S), for RSS operations, along with an estimation formula that establishes the quantitative relationship between system performance and configurations. Simulations validate that RC-S reduces average service time by 10.3% compared to the classical cooperative A* algorithm, while also improving throughput and runtime. Based on the performance analysis of RC-S, we develop a layout optimization model that considers system configurations, desired throughput, and costs to minimize expenses and determine the optimal layout. Numerical studies show that facility costs dominate at lower throughput levels, while labor costs prevail at higher throughput levels. Additionally, due to traffic efficiency limitations, RSS is well-suited for small-scale operations like end-of-supply-chain distribution centers.
Keywords: Logistics; Robotic sorting system; Robot management system; Performance evaluation; Layout design
@@ 1 Introduction
Improved logistics and delivery services have fueled the rapid growth of e-commerce in the 21st century. The development of stable supply chains, same-day or next-day delivery options, and hassle-free return policies increase consumer confidence in online shopping (BusinessWire,, 2023). In order to handle the increasing volume of online orders, a new generation of warehouses specifically catering to individual customers has become a hot topic for logistics companies. This kind of warehouse efficiently meets the demand for small orders with tight delivery schedules through the implementation of automated equipments (Boysen et al., 2019b, ). For instance, robotic mobile fulfill system (RMFS), which employs a rack-moving mechanism, is widely used in intra-warehouse logistics, such as Amazon’s KIVA systems (Wurman et al.,, 2008).
The main activities performed in warehouse include: (1) receiving, (2) transfer and put away, (3) order picking/selection, (4) accumulation/sorting, (5) cross-docking, and (6) shipping (De Koster et al.,, 2007). In this paper, we focus on the sorting process. Sorting involves categorizing and consolidating parcels according to their order information and shipping destinations. Conventional sorting systems commonly employ conveyor-based sorters , where the actuators move along with the conveyor belt, sequentially passing through each outbound station and releasing the loaded parcels at the appropriate locations (Boysen et al., 2019a, ). These systems are highly appreciated for its efficiency and stability, while suffering from the inflexibility and significant space occupation (Boysen et al.,, 2023).
@@
A novel sorting system utilizing mobile robots attracts increasing attention in recent years, which is called the robotic sorting system (RSS) (Zou et al.,, 2021). Figure 1(a) presents a dual-layer solution of RSS. The site is arranged with numerous delivery ports, each associated with a specific sorting category, gathering all the parcels of that category. A delivery port can be a bin, a cage cart, or a chute in a dual-layer structure. Robots deliver parcels to delivery ports, serving as the combination of conveyor belts and actuators in conventional sorting systems. Loading stations are located on the periphery of the site, where workers load parcels onto the robots. A top-mounted tray or conveyor actuator enables the robot to load or unload parcels. Figure 1(b) shows a specific model of sorting robot developed by Geekplus, a global technology company specialized in smart logistics and robots. The process for a robot to execute a sorting task is as follows:
@@
E-commerce order fluctuate during special selling seasons; therefore, warehouse throughput must be designed to accommodate these variations. Unlike conveyor-based sorting systems, the independent and modular nature of robotic operations contributes to the flexibility and scalability of RSS, making it well-suited to dynamic demands of e-commerce (Azadeh et al.,, 2017). During off-hours, warehouse managers can easily configure the control software to change the status of loading stations, the correspondences between outlets and destinations, and the aisle network topology (Xu et al.,, 2022). To leverage the flexibility, robotics companies introduce the innovative Robots-as-a-Service (RaaS) business model, allowing logistics clients to rent robots as needed. RaaS eliminates the high costs associated with purchasing and maintaining robots and peripheral equipment, enabling warehouse managers to adjust the number of rented robots based on demand fluctuations.
@@
Despite these advantages, a research gap remains in investigating the flexibility of RSS and minimizing system costs, as accurately modeling system efficiency is challenging. In robotic systems, robots rely on path planning and coordination algorithms to complete tasks. The problem of finding conflict-free trajectories for all robots is known as the multi-agent path finding (MAPF) problem. The MAPF problem has been proven to be NP-hard (Yu and LaValle,, 2013), which means that finding a relatively good solution can be time-consuming, especially when the scale of the robot system increases, resulting in the instability of RSS efficiency. Shi et al., (2021) analyzed a dataset from the China Post sortation center and found that the congestion effect significantly hampers robot efficiency in RSS, particularly when dealing with a substantial parcel flow. However, existing researches often overlook issues related to robot coordination, making the idealized assumption that there are no conflicts or deadlocks. Additionally, most of these studies tend to assume a fixed route travel time, excluding potential queuing delays on the road. As a result, their outcomes may become distorted, as the actual travel time of the robots deviates from the free-flow travel time. Furthermore, these studies do not adequately explore the flexibility of RSS, particularly the ability to adjust the number of robots.
@@
Given the existing gap, this paper aims to address critical challenges in designing an efficient RSS, focusing on three key areas. First, we seek to ensure the system operates both stably and efficiently in a setting where multiple robots collaborate, a non-trivial challenge given the complexities of robot interaction. Second, the development of a precise model to quantify system efficiency within the RSS is essential. Such a model will enable accurate throughput estimation, which is vital for understanding the relationship between the size of the facility and its sorting capacity. Third, we emphasize the need for management strategies that leverage the flexibility of RSS to achieve cost-effective responses to fluctuating demands. These targeted areas of improvement aim to enhance the overall performance and scalability of RSS implementations.
@@
To address the aforementioned challenges, we first explore the potential of centralized control strategy for sorting robots. To fully utilize global information while considering robot operational patterns and safety distance regulations, we adapt a network-level control strategy for coordinating autonomous vehicles and propose a collision-free spatio-temporal path planning method for RSS. An efficient heuristic algorithm is proposed to enhance the scalability. Second, building upon this control strategy, we derive an estimation formula for throughput in RSS given system configurations, which serves as an efficiency constraint in the planning problem. Finally, we propose an optimization model aimed at minimizing the initial investment and average operations costs under fluctuating demands. The solution includes layout design and resource allocation recommendations, specifically on how to adjust the number of robots and workers. Experiments provide insights into the cost structure of RSS under varying unit price conditions. To the best of our knowledge, this paper is among the first to incorporate the traffic issues of robots during operations into the configuration design of the RSS system. Our intention is to bridge operational and strategic planning in multi-robot coordination contexts, thereby contributing a comprehensive theoretical framework to support decision-making in such complex and flexible systems.
@@
The remainder of this paper is organized as follows. In section 2, we provide an extensive review of relevant research on RSS systems and conventional sorting systems, highlighting the contributions of this research. Section 3 offers a detailed problem description. In section 4, we introduce an innovative traffic management framework to coordinate multiple robots in a warehouse setting. Based on the proposed framework, section 5 presents a system throughput estimation formula, which provides quantitative measurement of layout configuration’s impact to system efficiency. In section 6, we demonstrate the superiority of our proposed traffic management framework compared with benchmarks, as well as validate the accuracy of throughput estimation formula. To emphasize the impact of traffic issues on RSS, experiments are conducted to show that queueing network model has estimation biases for throughput in certain scenarios. Moving on to section 7, we propose a layout optimization model to minimize the total costs and introduce an efficient solution approach. Section 8 presents a sensitivity analysis to investigate the optimal layout design and the corresponding cost structure under different unit cost, and insights distilled from the results. Finally, section 9 provides a summary of the entire paper.
@@ 2 Literature Review
The cost savings in labor through automation are particularly crucial in the picking process, and increased sorting efficiency enables parcel companies to offer convenient and low-cost same-day delivery services (Dekhne et al.,, 2019). However, the operations research community has scarcely explored the realm of robotized sorting systems, which represent the latest advance of warehouse automation. Among the existing literature, studies mainly focus on the efficiency of small parcel sorting systems in distribution centers. This body of work delves into the comprehensive assessment of transportation efficiency within the robot fleet under various system configurations. In the majority of these studies, robots’ sorting tasks are modeled as services in queuing network models, and the total throughput is the key performance evaluation metrics (Zou and Chen,, 2020; Zi and Gao, 2020a, ; Zi and Gao, 2020b, ; Zou et al.,, 2021; Xu et al.,, 2022). Zi and Gao, 2020a ; Zi and Gao, 2020b assumed that each robot moves independently. Zou and Chen, (2020) introduced a semi-open queuing network (SOQN) model and subsequently improved its theoretical aspects in their follow-up research (Zou et al.,, 2021), simplifying it to the closed queueing model (CQN) by assuming a sufficient high arrival rate of parcels. A queueing model takes into account the queuing behavior of robots at loading stations and outlets, partially addressing the issue of robot occupancy on the aisle. Furthermore, their second study compares how the minimum total costs of the system varied under different network topology and different unit cost of system components, given the target sorting efficiency. Xu et al., (2022) further investigated an RSS with a parcel-to-loading-station assignment mechanism, which could be regarded as a pre-sorting strategy. This study removes the constraint on the number of robots, resulting in an open queuing network (OQN) model. Experimental results demonstrates that the introduction of pre-sorting effectively reduces the average travel distance of the robots but might lead to extra congestion upstream at the same time.
Liu et al., 2019b and Tan et al., (2021) focused on the task assignment problem with the objective of minimizing the sorting makespan. Both adopts a travel time model (closed-form travel time expression, defined by Azadeh et al., (2017)), and the second research formulates a mixed-integer programming model. Compared to queuing network models, travel time models exhibit significant distortion when the number of robots is high, as they allow multiple robots to simultaneously occupy a facility. Boysen et al., (2023) conducted research on order sorting system, which involves additional tasks including Piece-to-Order assignment and Order-to-Collection Point assignment. Therefore, the primary focus of the referred study is not solely on scheduling and control of robots, and the travel time of robots to a specific outlet is assumed to be constant. Shi et al., (2021)’s study, on the other hand, investigates a human-robot hybrid sorting system, in which the scale of robot sorting is fixed, while the manual capacity is adjustable according to the demand. Regrettably, in this context, the considerable flexibility inherent in the robotic system has not been fully harnessed.
@@
Throughout the existing research, a limited number of studies consider the floor space cost while others relax the constraints on aisle resources. Furthermore, finding collision-free paths for multiple vehicles has been proved to be an NP-hard problem (Surynek,, 2010), and some state-of-the-art algorithms also struggle to completely eliminate the issue of road congestion resulting from an increasing number of robots. However, except the study by Shi et al., (2021), which employs a fitting method to predict the positive correlation between traffic flow and congestion, most of the studies neglect the mutual influence among multiple robots in aisle occupation for a compact system, leading to an misestimate of traffic capacity. Table 1 presents a comprehensive overview of the literature reviewed on the RSS, highlighting its relevance to our paper.
@@
Some researchers recognized the limitations of queuing models in capturing the impact of aisle layout design on multi-robot systems and chose to address the traffic issues by modeling them as MAPF problems. Wagner and Choset, (2011) proposed a novel M* algorithm that allowed for the planning of a larger number of robot trajectories. However, they faced challenges due to significant memory requirements and time complexity, and could not guarantee a feasible solution. Some researchers designed priority-based algorithms and integrated task assignment to reduce waiting time (Nguyen et al.,, 2019; Liu et al., 2019a, ), sacrificing solution optimality in exchange for stability and computational efficiency. Li et al., (2021) introduced an RHCR algorithm framework to tackle the windowed MAPF problem, providing an efficient solution for managing newly arrived orders in warehouse operations. However, despite their effectiveness, MAPF-based studies have faced challenges in deriving a closed-form expression for throughput estimation, which is crucial for creating reliable and efficient system configurations. Consequently, few studies on multi-robot systems comprehensively address both operational robot management and strategic layout design while considering their interplay.
In summary, existing research on RSS systems lacks sufficient focus on robot traffic management and overall system costs. This study addresses these gaps by first developing a dedicated robot traffic management method for RSS at the operational level. Building on this method, we establish an optimal layout and configuration design model aimed at minimizing overall system costs in the long run, ensuring that the results are practically applicable in real-world warehouses.